# Ageing-aware Energy Management for Residential Multi-Carrier Energy Systems

### v3

@@ 3.1 Power & thermal balances
The thermal balance comes in as:
@@ 3.2 Thermal modelling
The thermal assets are modelled in a linear way:
@@
and
@@
where $\eta$ denotes a conversion factor or efficiency, $Q_{\textrm{TESS}}$ is the capacity in kWh. To avoid imposing an arbitrary periodicity on the TESS, no terminal conditions are used on its $SoC_{\text{TESS},t}$. Thus, the TESS operates under its natural time constant.
@@ 3.3 Batteries
The remaining devices in the MCES are all battery-based ESS. Batteries have complex nonlinear dynamics, and several modeling techniques are presented in the literature [17]. In this work, models coming from empirical and physics-based approaches are used. The modeling is divided into two different sub-models: performance and ageing. Under the UMF, this is represented in the transition function $\tilde{S}^{M}_{b,t}(\tilde{S}_{b,t},x_{b,t}|\theta_{b,t})$, which contains both the perf. model $p_{b,t}^{M}(.)$ and the ageing model $d_{b,t}^{M}(.)$. The performance model predicts stored energy $SoC_{b,t}$ and terminal voltage $v_{t,b,t}$. The ageing model is used to update the parameters of $p^{M}_{b,t}(.)$, as shown in Fig. 3. Even though the change in parameters $\theta_{b,t}$ becomes significant after considerable ageing has occurred, optimizing it in the short term can lead to considerable savings in the long and medium term [46, 47, 15]. This is because in PB ageing models, the relationship between ageing states and control actions is explicit, and the policy can directly minimize it through the internal states $SoC_{b,t}$.
Figure 3: Storage asset transition function diagram $\tilde{S}^{M}_{b,t}$
@@
Finally, a terminal constraint is implemented to ease up feasibility and mitigate symmetries in the OCP of Eq. 8, as in:
$SoC_{\textrm{BESS},t^{\prime}_{1}}=SoC_{\textrm{BESS},t^{\prime}_{1}+24\text{hs}}$ (20)
where $t_{0}\leq t_{1}\leq H\leq T$. This way, the OCP is better computationally conditioned, but the planner still has the freedom to decide $SoC_{\textrm{BESS},24\textrm{hs}}$. The proposed terminal condition has two key properties: it is more flexible than fixing $SoC_{\text{BESS},24\text{hs}}=50\%$ , and it bounds the value function $V_{H}$ of the OCP. Ideally, no terminal condition would be used to freely use all 3 storage systems. Unfortunately, to solve such an unbounded OCP, an optimization horizon $H$ much larger than 48 hours would be required [48, 49].
@@ 3.3.1 Performance models $p_{b,t}^{M}$
For the performance submodel, two alternatives have been implemented: a simple bucket model (BM) and a first-order equivalent circuit model (ECM). A basic BM of the operation of a battery assumes that its output voltage $v_{t}$ is linear with the state of charge $SoC$, assuming no voltage drop. Hence, the only equations of this model are:
@@
where $i_{b,t}$ is the current passing through the cell, $OCV_{b,t}$ is the open circuit voltage, $\eta_{c}$ is the Coulombic efficiency [18] , and $Q_{b,t}$ is the cell capacity in Ah. Each battery pack is assumed to be organized as a series connected module (SCM) where $N_{s/p,\ b}$ are the series cells per branch and parallel branches, respectively. In this model, the most relevant parameter in $\theta_{b}$ is the $Q_{b}$.
@@
where $i_{R_{1},b,t}$ is the current flowing through $R_{1}$ in Fig. 4. Eqs. 21a, 21b and 21c are maintained. The ECM incorporates the series voltage drop that limits power output and the first-order diffusion dynamics. Here, the relevant parameters are $\theta=[Q,R_{0}]^{T}$ that usually define the cell’s state of health $SoH$.
@@
For the ageing models, the first alternative is an empirical sub-model presented by [32]. The empirical sub-model reduces all the degradation mechanisms into calendar and cyclic ageing.
@@
where the fitting parameters $c_{1,...,5}$ are taken from [50, 10, 32] and $t_{0,b}$ is the elapsed lifetime of the battery $b$.
For the physics-based alternative, the reduced order model (PBROM) from [15] is used. It accounts for two degradation mechanisms: the solid electrolyte interface (SEI) and active material loss (AM). The author also presents a PBROM for Li-plating, but given the low C-rate and standard temperature range of this application, it will not be included.
@@
where $\eta_{k}$ is the SEI side reaction kinetic overpotential, $z$ is the Li stochiometry of the cell, $OCV_{n}$ is the open-circuit voltage of the anode made with an empirical fit, $OCV_{s}=0.4V$ is the side reaction open-circuit voltage, and $T$ is the cell temperature. It is assumed that the temperature $T$ is constant over time and is controlled by the local primary control system. For the BESS this assumption could be challenged given that the BESS is owned/operated by the same house owners. In the case of the EV, the ownership of the car battery pack sometimes corresponds to the OEM/car manufacturer, limiting access and control over it. The rest of the parameters can be found in the A.
@@
which is later used again in 25.
@@ 3.4 Electric Vehicle
From the point of view of a residential building, the EV s are a BESS with availability constraints and certain requirements regarding their $SoC$ at departure time $t_{\text{dep}}$. For the availability $\gamma$, the probability distributions of departure ($t_{\text{dep}}$) and arrival ($t_{\text{arr}}$) times can be described as random variables $t_{\textrm{dep/arr}}\sim\mathcal{T_{\textrm{dep/arr}}}$, whose distributions $\mathcal{T}_{\text{dep/arr}}$ are taken from Elaad [51]. The availability $\gamma_{t}$ will then be:
@@
Summing up, the three different storage systems have three different ways of handling their terminal set and value function bounds. The BESS has a periodic condition in Eq. 20 , which is a flexible alternative to the classic terminal daily periodicity. The EV follows a soft-tracking problem at $t_{\text{dep}}$. The TESS has no terminal condition. This flexible design is the key contribution to integrating all the storage systems and the non-convex PB models to OCP in Eq. 8.
@@
1:Define setpoint sequences $\mathcal{P}_{a,[t,t+24\text{hs}]}$
2:Define exogenous information sequences $\mathcal{W}_{[t+1,t+1+24\text{hs}]}$
3:Recalculate $P_{\textrm{TESS},t}$ and $P_{\textrm{grid},t}$ with Eqs. 13 , 14, $\mathcal{P}_{a,[t,t+24\text{hs}]}$, and $\mathcal{W}_{a,[t+1,t+1+24\text{hs}]}$
4:for $t\in\mathcal{D}_{t}$ do
5:  Simulate $SoC_{\textrm{TESS},t}$ using Eq. 17
6:  Simulate $b$ performance using $p_{b}^{M}(.)$ PBROM [52]
7:  Simulate $b$ degradation using $d_{b,t}^{M}(.)$ PBROM [15]
8:end for
9:Feed-back $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$ to the planner
@@
The whole process is defined in Algorithm 1. First, the power setpoints must be adjusted for the grid and TESS because the forecast used in $X_{t}^{\pi}$ will never be the same as the actual exogenous inputs. Take a look at the balances, Eq. 13 and 14, which contain the loads and solar generation. It is clear that $\tilde{P}\neq P$ and a device must compensate for that difference. Thus, the simulator $S_{a,t}^{M}$ recalculates:
@@
Second, once these powers have been adjusted, the simulator uses these powers to obtain the true/actual/fidelity state sequence $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$. For the TESS it recalculates Eq. 17. For the $b$, it uses LiiBRA.jl [52] to swiftly simulate PBROM s of the performance of the battery [17, 53]. After that, the models from Jin [15] are used to calculate the true degradation outcome of the decisions $P_{a,t}^{*}$. Again, the reader must remember that the capacity fade (decrease in $Q_{b,t}$) is modeled in both the simulator $S^{M}_{a,t}$ and the approximate model of the planner $\tilde{S}_{a,t}^{M}$, whereas the power fade (increase in $R_{0,as,t}$) is only addressed in the simulator $S^{M}_{a,t}$. Finally, if an action $P_{b,t}^{*}$ causes a future state to go out of bounds ($S_{b,t+1}\leq\underline{S}_{b}$ or $S_{b,t+1}\geq\overline{S}_{b}$), the remaining actions are rejected and the $b$ remains in that state (either $\overline{S}_{b}$ or $\underline{S}_{b}$) until the next day. Finally, the carriers are rebalanced if necessary.
The final state sequence $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$ is then fed back to the optimization-based planner. For practical implementation, in which the simulator is, in fact, an experimental setup, an online state observer is necessary to feed back the states to the EMS. This is particularly important for the ESS [17, 18, 54].
@@ 5 Case studies
1:Initialize hyperparameters $t_{0},\ \Delta t,\ t_{W},\ w,\ n_{d}$
@@
4:  Solve the deterministic OCP, Eq. 8, and obtain $P_{a,t}^{*}$.
5:  Simulate $S_{a,t+1}=S_{a,t}^{M}(S_{a,t},P^{*}_{a,t},W_{t+1})\ \forall t\in\mathcal{D}_{t}=[t;t+24\text{hs}]$, using Algorithm 1
6:  Update forecasts in $\tilde{B}_{a,t}$
7:  Move time window $\mathcal{D}^{\pi}_{t}\leftarrow\mathcal{D}^{\pi}_{t}+24\text{hs}$;
8:end for
The building has a grid connection with a smart meter with 15min resolution. The grid power $P_{\textrm{grid}}$ is included in the state vector $S_{a,t}$.
The system is composed of a 5kWp SPV, a 20kWh BESS with nickel manganese cobalt oxides (NMC) or Lithium iron phosphate (LFP) cells, two 12.5kW EV charging points, a 4kWe heat pump, a 2.7kWth solar thermal collector, a 200kWh TESS, a 6kWp electrical load, a 1kWp thermal demand, and 10kW LV grid connection. Power consumption profiles ($P_{\textrm{load}}^{\textrm{e}}$) were constructed for a year using data from 2021 to 2023 from TU Delft’s Green Village smart meter data [55]. The output of the SPV is taken from [56, 57], the market prices $\lambda$ are taken from the EPEX day-ahead auction, and $\lambda_{\textrm{buy}}=0.95\lambda_{\textrm{sell}}$ [58], and the heat demand $P_{\textrm{load}}^{\textrm{th}}$ was modeled as [34].
The cells used are SANYO NCR18650 cells for NMC as in [15] and A123 cells for LFP [31]. Their datasets were taken from PyBaMM [59] and LiiBRA [52]. The ECM s for both cell types were fitted from synthetic cells simulated in PyBaMM, following the methodology in [17]. Once the simulated profiles are ready, the ECM parameters can be identified using subspace system identification as in [60]. The parameters can be found in A, Tables 4 and 5. The capacity fade cost is assumed to be $c_{\text{loss}}=1.2$ €/Ah, roughly 280-310€/kWh, depending on the average voltage.
Figure 6: Monthly dispatch of MCES under CPBDeg-DA planner for summer (left) and winter (right).
The OCP and simulations were modelled and run using Julia [61], JuMP [62], and InfiniteOpt [63]. The chosen solver was KNITRO’s active sets algorithm from Artelys since it can handle NLP with complementarity constraints [64]. All simulations were run using an Intel CPU at 2.60GHz and 4 processors and 32GB of RAM.
5.1 Case Study I: Day-ahead planning
To test and validate our EMA the planner was simulated for two standard months (summer and winter) using 2023 data from the previously mentioned sources. Both are presented in Fig. 5. To quantify the impact of each performance and ageing model 3 day-ahead planners, 2 benchmarks and our proposed approach, were implemented:

### v4

@@ 3.1 Power & thermal balances
Assuming an indoor temperature setpoint decided either by the user or the thermal system [Alpizar-Castillo2024ModellingHouse, Damianakis2023Risk-averseConsumption, Slaifstein2026SequentialMPC], the building’s thermal balance comes in as:
@@ 3.2 Thermal modelling
The thermal assets are modelled in a linear way. The building has an air-to-water heat-pump that transforms electric power $P_{\text{HP},t}^{\text{e}}$ to heat-flow $P_{\text{HP},t}^{\text{th}}$ with a coefficient of performance $\eta_{\text{HP}}$:
@@
where $\eta_{\text{HP}}$ is assumed constant due to low variation of it during the year [Alpizar-Castillo2024ModellingHouse].
The thermal solar collector converts irradiance into heat, and assuming a constant linear relationship with the generated SPV power $P_{\text{PV},t}$ as:
@@
This work considers an underground, perfectly mixed water tank with independent charge and discharge coils as a TESS. Assuming no mass exchange between the TESS and the piping system the TESS the state of charge is:
@@
where $\eta_{\text{TESS}}$ denotes the heat-exchanger efficiency, $Q_{\textrm{TESS}}$ is the capacity in kWh. To avoid imposing an arbitrary periodicity on the TESS, no terminal conditions are used on its $SoC_{\text{TESS},t}$. Thus, the TESS operates under its natural time constant.
@@ 3.3 Batteries
The remaining devices in the MCES are all battery-based ESS. Batteries have complex nonlinear dynamics, and several modeling techniques are presented in the literature [Plett2015]. In this work, models coming from empirical and physics-based approaches are used. The modeling is divided into two different sub-models: performance and ageing. Under the UMF, this is represented in the transition function $\tilde{S}^{M}_{b,t}(\tilde{S}_{b,t},x_{b,t}|\theta_{b,t})$, which contains both the perf. model $p_{b,t}^{M}(.)$ and the ageing model $d_{b,t}^{M}(.)$. The performance model predicts stored energy $SoC_{b,t}$ and terminal voltage $v_{t,b,t}$. The ageing model is used to update the parameters of $p^{M}_{b,t}(.)$, as shown in Fig. 3. Even though the change in parameters $\theta_{b,t}$ becomes significant after considerable ageing has occurred, optimizing it in the short term can lead to considerable savings in the long and medium term [Cao2020MultiscaleModels, Movahedi2024ExtraMechanism, Jin2022, Reniers2021]. This is because in PB ageing models, the relationship between ageing states and control actions is explicit, and the policy can directly minimize it through the internal states.
Figure 3: Battery storage asset transition function diagram $\tilde{S}^{M}_{b,t}$
@@
Finally, assuming that there is a strong daily usage pattern a terminal constraint is implemented to ease up feasibility and mitigate symmetries in the optimal control problem (OCP) of Eq. 8. Following [Slaifstein2026SequentialMPC], the daily periodical constraint in the first optimization of the day ($X^{\text{DA}}_{t}$ in Fig. 2) is:
$\tilde{SoC}_{\textrm{BESS},tt^{\prime}_{1}}^{\textrm{DA}}=\tilde{SoC}_{\textrm{BESS},tt^{\prime}_{1}+24hs}^{\textrm{DA}}$ (18)
and in all other steps $X^{\text{MPC}}_{t}$ the periodic condition is:
$\tilde{SoC}_{\textrm{BESS},tt^{\prime}_{0}}^{\textrm{MPC}}=\tilde{SoC}_{\textrm{BESS},tt^{\prime}_{0}+H^{\textrm{MPC}}}^{\textrm{MPC}}$ (19)
where $t_{0}\leq t_{1}\leq H\leq T$. This way, the OCP is better computationally conditioned, but the planner still has the freedom to decide the state-of-charge at the begining of each day ($SoC_{\textrm{BESS},24\textrm{hs}}$). The proposed terminal condition has two key properties: it is more flexible than fixing $SoC_{\text{BESS},24\text{hs}}=50\%$ , and it bounds the value function $V_{H}$ of the OCP. Ideally, no terminal condition would be used to freely use all 3 storage systems. Unfortunately, to solve such an unbounded OCP, an optimization horizon $H$ much larger than 48 hours would be required [Grune2017NonlinearAlgorithms, Prat2024HowProblems].
@@ 3.3.1 Performance models $p_{b,t}^{M}$
For the performance submodel, two alternatives have been implemented: a simple Coulomb counting or bucket model (BM) and a first-order equivalent circuit model (ECM). A basic BM of the operation of a battery assumes that its output voltage $v_{t}$ is linear with the state of charge $SoC$, assuming no internal resistance, and no-diffusion dynamics. Hence, the only equations of this model are:
@@
where $i_{b,t}$ is the current passing through the cell, $OCV_{b,t}$ is the open circuit voltage, $\eta_{c}$ is the Coulombic efficiency [Plett2016] , and $Q_{b,t}$ is the cell capacity in Ah. Each battery pack is assumed to be organized as a series connected module (SCM) where $N_{s/p,\ b}$ are the series cells per branch and parallel branches, respectively. In this model, the most relevant parameter in $\theta_{b}$ is the $Q_{b}$.
@@
where $i_{R_{1},b,t}$ is the current flowing through $R_{1}$ in Fig. 4. Eqs. 20a, 20b and 20c are maintained. The ECM incorporates the series voltage drop that limits power output and the first-order diffusion dynamics. Here, the relevant parameters are $\theta=[Q,R_{0}]^{T}$ that usually define the cell’s state of health $SoH$.
@@
For the ageing models, the first alternative is an empirical sub-model presented by [Wang2014]. The empirical sub-model reduces all the degradation mechanisms into calendar and cyclic ageing.
@@
where the fitting parameters $c_{1,...,5}$ are taken from [SlaifsteinIECON23, Vermeer2022b, Wang2014] and $t_{0,b}$ is the elapsed lifetime of the battery $b$.
For the physics-based alternative, the reduced order model (PBROM) from [Jin2022] is used. It accounts for two degradation mechanisms: the solid electrolyte interface (SEI) and active material loss (AM). The author also presents a PBROM for Li-plating, but given the low C-rate and standard temperature range of this application, it will not be included.
@@
where $\eta_{k}$ is the SEI side reaction kinetic overpotential, $z$ is the Li stochiometry of the cell, $OCV_{n}$ is the open-circuit voltage of the anode made with an empirical fit, $OCV_{s}=0.4V$ is the side reaction open-circuit voltage, and $T$ is the cell temperature. It is assumed that the temperature $T$ is constant over time and is controlled by the local primary control system of the BESS and EV. The rest of the parameters can be found in the A.
@@
which is later used again in 24.
@@ 3.4 Electric Vehicle
From the point of view of a residential building, the EV s are a BESS with availability constraints and certain requirements regarding their $SoC$ at departure time $t_{\text{dep}}$. For the availability $\gamma$, the probability distributions of departure ($t_{\text{dep}}$) and arrival ($t_{\text{arr}}$) times can be described as random variables $t_{\textrm{dep/arr}}\sim\mathcal{T_{\textrm{dep/arr}}}$, whose distributions $\mathcal{T}_{\text{dep/arr}}$ are taken from Elaad [elaad]. The availability $\gamma_{t}$ will then be:
@@
Summing up, the three different storage systems have three different ways of handling their terminal set and value function bounds. The BESS has a periodic conditions Eqs. 18 and 19 , which is a flexible alternative to the classic terminal daily periodicity. The EV follows a soft-tracking problem at $t_{\text{dep}}$. The TESS has no terminal condition. This flexible design is the key contribution to integrating all the storage systems and the non-convex PB models to the eMPC of Eq. 8.
@@
1:Define setpoint $P_{a,t}$
2:Define exogenous information $W_{t+1}$
3:Recalculate $P_{\textrm{TESS},t}$ and $P_{\textrm{grid},t}$ using Eqs. 11 & 12, with $P_{a,t}$, and $W_{t+1}$
4:Simulate $SoC_{\textrm{TESS},t}$ using Eq. 15
5:Simulate $b$ performance using $p_{b}^{M}(.)$ PBROM [Planden2022]
6:Simulate $b$ degradation using $d_{b,t}^{M}(.)$ PBROM [Jin2022]
7:Feed-back $S_{a,t}$ to the planner
@@
The whole process is defined in Algorithm 1. First, the power setpoints must be adjusted for the grid and TESS because the forecast used in $X_{t}^{\pi}$ will never be the same as the actual exogenous inputs. Take a look at the balances, Eq. 11 and 12, which contain the loads and solar generation. It is clear that $\tilde{P}\neq P$ and a device must compensate for that difference. Thus, the simulator $S_{a,t}^{M}$ recalculates:
@@
Second, once these powers have been adjusted, the simulator uses these powers to obtain the true/actual/fidelity state $S_{a,t}$. For the TESS it recalculates Eq. 15. For the $b$, it uses LiiBRA.jl [Planden2022] to swiftly simulate PBROM s of the performance of the battery [Plett2015, Rodriguez2019]. After that, the models from Jin [Jin2022] are used to calculate the true degradation outcome of the decisions $P_{a,t}^{*}$. Again, the reader must remember that the capacity fade (decrease in $Q_{b,t}$) is modeled in both the simulator $S^{M}_{a,t}$ and the approximate model of the planner $\tilde{S}_{a,t}^{M}$, whereas the power fade (increase in $R_{0,b,t}$) is only addressed in the simulator $S^{M}_{a,t}$. Finally, if an action $P_{b,t}^{*}$ causes a future state to go out of bounds ($S_{b,t+1}\leq\underline{S}_{b}$ or $S_{b,t+1}\geq\overline{S}_{b}$), the action is rejected and modified leaving $b$ on the bound (either $\overline{S}_{b}$ or $\underline{S}_{b}$) until the next $t$. Finally, the carriers are rebalanced if necessary.
The final state $S_{a,t}$ is then fed back to the optimization-based controller. For practical implementation, in which the simulator is, in fact, an experimental setup, an online state observer is necessary to feed back the states to the EMS. This is particularly important for the ESS [Plett2015, Plett2016, Plett2024BatteryMethods, Fan2023NondestructivePaths].
(a) Simulated and predicted SoCSoC for the EV for T=24T=24hs and open-loop simulation.
(b) Full simulation and reported down sampled state.
Figure 5:
Despite the inherent inaccuracies of the simplified models used within the EMS planners $\tilde{S}_{a,t}^{M}(\cdot)$, the informed $SoH_{b,t}$ remains accurate due to the high-fidelity feedback provided by $S_{b,t}^{M}(\cdot)$. It is assumed that the primary control of the storage $b$ will maintain the current spikes within tight bounds, avoiding any consequent voltage deviations within the sampling time $\Delta t=15$min of the EMS. Thus, since the simulator $S_{b,t}^{M}(\cdot)$ operates at a $\Delta t_{s}=1$s resolution, the informed degradation captures the true physical response of the battery, including peak state values. These are later down-sampled to the $\Delta t=15$min resolution of the planner $X^{\pi}$. The process effectively reduces the planner’s error to a mere 15-minute communication lag (Fig. 5) and a zero-order-hold approximation. By utilizing the realized trajectories of $SoC_{b,t}$ and $v_{t,b,t}$ from the fidelity simulator $S_{a,t}^{M}(\cdot)$ rather than the planner’s internal estimates, the system ensures that the reported capacity fade and degradation states reflect the actual operational stress on the battery, regardless of the planner’s complexity.
@@ 5 Case studies
1:Initialize hyperparameters $t_{0},\ \Delta t,\ H^{\text{DA/MPC}},\ w,\ n_{d}$
@@
4:  Solve the day-ahead policy $X^{\text{DA}_{t}}$ and obtain schedule $\mathcal{P}_{a,[t,H^{\text{DA}}]}^{\textrm{DA}}$.
5:  for $t\in 0:H^{\text{MPC}}$ do
6:   Solve the deterministic continuous time policy $X^{\text{MPC}_{t}}$ and obtain action $P_{a,t}^{\textrm{MPC}}$.
7:   Simulate $S_{a,t+1}=S_{a,t}^{M}(S_{a,t},P^{\textrm{MPC}}_{a,t},W_{t+1})\$
8:   Update forecasts in $B_{a,tt^{\prime}}^{\text{MPC}}$
9:   Move time window $t\leftarrow t+\Delta t^{\text{MPC}}$;
10:  end for
11:end for
The building has a grid connection with a smart meter with 15min resolution. The grid power $P_{\textrm{grid}}$ is included in the state vector $S_{a,t}$. The system is composed of a 4kWp SPV, a 15.6kWh/12.5kW BESS with nickel manganese cobalt oxides (NMC) or Lithium iron phosphate (LFP) cells, one 55.6kWh/12.5kW EV charging points, a 4kWe heat pump, a 2.7kWth solar thermal collector, a 200kWh TESS, a 2.5kWp electrical load, a 1.5kWp thermal demand, and 17kW LV grid connection. Power consumption profiles ($P_{\textrm{load}}^{\textrm{e}}$) were constructed for a year using data from 2021 to 2023 from TU Delft’s Green Village smart meter data [GreenVillage]. The output of the SPV is taken from [Smets2016, Diab2023], the market prices $\lambda$ are taken from the EPEX day-ahead auction, and $\lambda_{\textrm{buy}}=0.95\lambda_{\textrm{sell}}$ [2023EPEXServices], and the heat demand $P_{\textrm{load}}^{\textrm{th}}$ was modeled as [Alpizar-Castillo2024ModellingHouse].
The cells used are SANYO NCR18650 cells for NMC as in [Jin2022] and A123 cells for LFP [Prada2013Simulations]. Their datasets were taken from PyBaMM [Sulzer2021PythonPyBaMM] and LiiBRA [Planden2022]. The ECM s for both cell types were fitted from synthetic cells simulated in PyBaMM, following the methodology in [Plett2015]. Once the simulated profiles are ready, the ECM parameters can be identified using subspace system identification as in [Plett2004b]. The parameters can be found in A, Tables 7 and 8 along with a more detailed explanation of the identification pipeline. The capacity fade cost is assumed to be $c_{\text{loss}}=1.2$ €/Ah, roughly 280-310€/kWh, depending on the average voltage.
The simulations were modelled and run using Julia [julia], JuMP [JuMP], and InfiniteOpt [InfiniteOpt]. The chosen solver was KNITRO from Artelys [KNITRO]. All simulations were run using an Intel CPU at 2.60GHz, 4 processors, and 32GB of RAM. The base library for building and simulating EMS policies is open-source and free to use in EMSmodule [Slaifstein2026SequentialMPC]. Since this is a non-linear non-convex OCP, global optimality can not be generally guaranteed [Boyd2011ConvexOptimization, Grune2017NonlinearAlgorithms]. Having this in mind, the implementation of the EMS approaches this limitation by: (i) using warm-starts as initial guess to ensure convergence to the same local optima, (ii) using the KNITRO tuner and multi-start in each individual step to increase the chance of finding the best local optima [KNITRO].
5.1 Case Study I: Degradation model comparison
To test and validate our EMA the closed-loop control was simulated for two standard months (summer and winter) using 2023 data from the previously mentioned sources. Both are presented in Fig. 6. To quantify the impact of each performance and ageing model 3 eMPC controllers, 2 benchmarks and our proposed approach, were implemented: